• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 16:35:41 on November 02, 2009

    Entry number 297182

    keyword=R-HRS S1 and S2 Status (as is)

    Figures 1-4 show the S1 and S2 right HRS ADC and TDC spectra as they are
    now. I have made no changes to the HVs.

    For S1, the third and fourth left-side signals are missing. The TDC
    signals are there for these channels. All S2 ADC's and TDC's have a
    signal. I will work on evaluating the noise in the TDC spectra for both

    A copy of this log entry has been emailed to: rsubedi,bogdanw,xiaochao,reimer

    R-HRS S1 ADC

    R-HRS S1 TDC

    R-HRS S2 ADC

    R-HRS S2 TDC